When originally purchased, a vehicle costing $23,940 had an estimated useful life of 8 years and an estimated salvage value of $2,100. After 4 years of straight-line depreciation, the asset's total estimated useful life was revised from 8 years to 6 years and there was no change in the estimated salvage value. The depreciation expense in year 5 equals:
Answer #1

Answer : $6,510

Depreciation = (cost - Salvage Value) / Life in YEars

Depreciation = (23,940-2100) /8 = $2730

Book Value for the start of the first year = 23940-(2730*4) = $13,020

Depreicaion for the 5th year = 13,020/2 =$6,510(Answer)